OpenStudy (anonymous):

If the points (–2, 2), (–4, 4), (2, –2), and (4, –4) are joined to form a straight line, at what point does the line intersect the y-axis? A. (0, 0) B. (4, –4) C. (2, 0) D. (0, –2)

OpenStudy (alexandervonhumboldt2):

well the easiest way is to make a equation: y=mx+b , where m is slope and is y-intercept first we need to find slope: the slope formula is \[\frac{ y_1-y_2 }{ x_1-x_2 }\] now substitute 2 of your points: \[\frac{ 2-4 }{ -2+4 }=\frac{ -2 }{ 2 }=-1\] so the slope is -1. m=-1. y=-x+b Now substitute one of your coordinates into the equation; (-2, 2) 2=-(-2)+b b=2-2 b=0 b=0, b is y-intercept, thus the y-intercept is 0.

OpenStudy (alexandervonhumboldt2):

here are the steps for you in example of how to find an equation of any line: Steps for finding the equation in example. For example, write an equation of line with points (1, 3) and (-8, 6) y=mx+b, where m is slope and b is y-intercept. First of all we need to find slope (m). To find it, we need to use the slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 } ~~ or~~ \frac{ y_2-y_1 }{ x_2-x_1 }\] Substitute our points: \[\frac{ 3-6 }{ 1+8 }=\frac{ -3 }{ 9 }=-\frac{ 1 }{ 3 }\]. Thus our slope is -1/3. Now we have that m=-1/3. Substitute: y=-1/3x+b. Now we need to find the y-intercept of out equation. To do this we need to substitute one of our points into the equation and solve: 3=-1/3*1+b b=3+1/3=10/3 Thus our equation is \[y=-\frac{ 1 }{ 3 }x+\frac{ 10 }{ 3 }\].

OpenStudy (anonymous):

Wow youre awesome. Thank you so much...=.)